# Rational solutions of Painlevé V from Hankel determinants and the asymptotics of their pole locations

### v2

@@ 2.2 From orthogonal polynomials to Lax Pairs
$\oint_{\gamma}\frac{\widetilde{p}_{n-1}(w){\rm e}^{\theta(w)}{\mathrm{d}}w}{2i\pi\,w}=-\dfrac{\begin{vmatrix}\mu_{0}&\mu_{1}&\mu_{2}&\dots&\mu_{n-1}\\ \mu_{1}&\mu_{2}&\mu_{3}&\dots&\mu_{n}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ \mu_{n-2}&\mu_{n-1}&\mu_{n}&\dots&\mu_{2n-3}\\ \mu_{-1}&\mu_{0}&\mu_{1}&\dots&\mu_{n-2}\end{vmatrix}}{\tau_{n}}=(-1)^{n}\dfrac{\begin{vmatrix}\mu_{-1}&\mu_{0}&\mu_{1}&\dots&\mu_{n-2}\\ \mu_{0}&\mu_{1}&\mu_{2}&\dots&\mu_{n-1}\\ \mu_{1}&\mu_{2}&\mu_{3}&\dots&\mu_{n}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ \mu_{n-2}&\mu_{n-1}&\mu_{n}&\dots&\mu_{2n-3}\end{vmatrix}}{\tau_{n}}$ (2.47)
@@
$\oint_{\gamma}\frac{p_{n}(w){\rm e}^{\theta(w)}{\mathrm{d}}w}{2i\pi\,w}=\dfrac{1}{2i\pi}\dfrac{\begin{vmatrix}\mu_{0}&\mu_{1}&\mu_{2}&\dots&\mu_{n}\\ \mu_{1}&\mu_{2}&\mu_{3}&\dots&\mu_{n+1}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ \mu_{n-1}&\mu_{n}&\mu_{n+1}&\dots&\mu_{2n-1}\\ \mu_{-1}&\mu_{0}&\mu_{1}&\dots&\mu_{n-1}\end{vmatrix}}{\tau_{n}}=\dfrac{(-1)^{n}}{2i\pi}\dfrac{\begin{vmatrix}\mu_{-1}&\mu_{0}&\mu_{1}&\dots&\mu_{n-1}\\ \mu_{0}&\mu_{1}&\mu_{2}&\dots&\mu_{n}\\ \mu_{1}&\mu_{2}&\mu_{3}&\dots&\mu_{n+1}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ \mu_{n-1}&\mu_{n}&\mu_{n+1}&\dots&\mu_{2n-1}\end{vmatrix}}{\tau_{n}}.$ (2.48)
@@
Inserting (2.45), (2.46), (2.47), (2.48) into (2.44), simplifying and rearranging the rows of the resulting determinants we obtain
@@
$(Y_{n}(0)^{-1}Y_{n}^{\prime}(0))_{11}=-\dfrac{\begin{vmatrix}\mu_{0}&\mu_{2}&\dots&\mu_{n}\\ \mu_{1}&\mu_{3}&\dots&\mu_{n+1}\\ \vdots&\vdots&\ddots&\vdots\\ \mu_{n-1}&\mu_{n+1}&\dots&\mu_{2n-1}\end{vmatrix}}{\tau_{n}}\dfrac{\begin{vmatrix}\mu_{-1}&\mu_{0}&\mu_{1}&\dots&\mu_{n-2}\\ \mu_{0}&\mu_{1}&\mu_{2}&\dots&\mu_{n-1}\\ \mu_{1}&\mu_{2}&\mu_{3}&\dots&\mu_{n}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ \mu_{n-2}&\mu_{n-1}&\mu_{n}&\dots&\mu_{2n-3}\end{vmatrix}}{\tau_{n}}+$
$+\dfrac{\begin{vmatrix}\mu_{0}&\mu_{2}&\dots&\mu_{n-1}\\ \mu_{1}&\mu_{3}&\dots&\mu_{n}\\ \vdots&\vdots&\ddots&\vdots\\ \mu_{n-2}&\mu_{n}&\dots&\mu_{2n-3}\end{vmatrix}}{\tau_{n}}\dfrac{\begin{vmatrix}\mu_{-1}&\mu_{0}&\mu_{1}&\dots&\mu_{n-1}\\ \mu_{0}&\mu_{1}&\mu_{2}&\dots&\mu_{n}\\ \mu_{1}&\mu_{2}&\mu_{3}&\dots&\mu_{n+1}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ \mu_{n-1}&\mu_{n}&\mu_{n+1}&\dots&\mu_{2n-1}\end{vmatrix}}{\tau_{n}}.$ (2.49)
@@
We now use the Desnanot-Jacobi identity, which can be stated as follows; given a square matrix $M$ denote by $M^{[a_{1},\dots,a_{r}][b_{1},\dots,b_{r}]}$ the matrix obtained by deleting the rows $a_{1},\dots,a_{r}$ and columns $b_{1},\dots,b_{r}$; then
@@
$M:=\begin{vmatrix}\mu_{-1}&\mu_{0}&\mu_{1}&\dots&\mu_{n-1}\\ \mu_{0}&\mu_{1}&\mu_{2}&\dots&\mu_{n}\\ \mu_{1}&\mu_{2}&\mu_{3}&\dots&\mu_{n+1}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ \mu_{n-1}&\mu_{n}&\mu_{n+1}&\dots&\mu_{2n-1}\end{vmatrix}\in{\rm gl}(n+1).$ (2.51)
@@
$\partial_{t}\mu_{j}(t)=\oint_{\gamma}z^{j-1}\left(1-\frac{1}{z}\right)^{{\rho}}z^{{K}}{\rm e}^{\frac{t}{z}}{\mathrm{d}}z=\mu_{j-1}(t),\ \ \ \ j\in{\mathbb{Z}}.$ (2.53)
This allows us to show that
$\partial_{t}\det\tau_{n}=\det M^{[n+1][2]}$ (2.54)
and hence
$(Y_{n}(0)^{-1}Y_{n}^{\prime}(0))_{11}=-\partial_{t}\ln\tau_{n}(t).$ (2.55)
This concludes the proof. $\blacksquare$
@@ 3 Asymptotic analysis
The goal of this second part of the paper is to study the behaviour of the poles of the rational solution or, which is the same, the zeros of the Hankel determinant $\tau_{n}(t)$ as $n\to\infty$. More precisely, and in line with similar investigations done for rational solution of PII [17, 6, 18] we are going to re-scale the zeros concurrently with $n$ and study the behaviour of the zeros in the $s=\frac{t}{n}$–plane. Namely we study the sequence of functions (polynomials)
$\mathcal{T}_{n}(s):=\tau_{n}(ns;{K},{{\rho}})$ (3.1)
@@
$\tau_{n}(t;{K},{{\rho}}):=\det\left[\oint_{\gamma}z^{a+b+{K}}\left(1-\frac{1}{z}\right)^{{\rho}}{\rm e}^{\frac{t}{z}}\frac{{\mathrm{d}}z}{2i\pi}\right]_{a,b=0}^{n-1}.$ (3.2)
We emphasize that we are not scaling the parameters ${{\rho}}\in{\mathbb{C}}\setminus{\mathbb{Z}}$ nor $K\in{\mathbb{Z}}$ as $n\to\infty$; a different asymptotic analysis would be needed if we were to set ${{\rho}}\propto n$ and/or $K\propto n$. We will briefly point out the main differences when appropriate.
The Theorem 2.1 can be re-formulated for convenience in the following form:
Theorem 3.1
The function $\mathcal{T}_{n}(s):=\tau_{n}(ns;K,{{\rho}})$ (3.1) is zero if and only if the RHP 2.1 has no solution (with $t=ns$).
The asymptotic analysis of RHP 2.1 with $t=ns$ falls within the purview of the Deift–Zhou asymptotic method that was brought to fruition in [26] for orthogonal polynomials, see also [11, 25].
@@
  1. 1.
    Construct a “$g$-function”; this is a function $g(z;s)$ whose domain of analyticity $\mathscr{D}_{s}$ depends (continuously) on the parameter $s$. It can be expressed as an Abelian integral on a hyperelliptic Riemann surface $\mathcal{R}$ whose genus depends on $s$. The $g$–function is uniquely characterized by a free-boundary value problem and certain inequalities which will be described in Sec. 4.
  2. 2.
    Use the constructed $g$–function to normalize RHP 2.1 and express a new matrix-valued function $W(z)$ which satisfies a new RHP.
  3. 3.
@@
The key concept is that the solvability of the initial RHP 2.1 is equivalent to the solvability of its final reformulation hinted at in the list above. Thus, if we can guarantee that the final reformulation is solvable for $s$ in a suitable domain, we can conclude that there are no zeros of $\mathcal{T}_{n}(s)$ in that domain (at least for $n$ sufficiently large).
Let us give a visual overview of the result; in Fig. 1 we can see several instances of plots of the zeros of $\mathcal{T}_{n}(s)$ for various values of ${{\rho}}$; the common feature, which will be proved, is that all the zeros lie, asymptotically for large $n$, within the “Eye of the Tiger” region ($EoT$), marked by the black arcs; the equation of these black arcs is given implicitly in (4.26).
Also indicated is a grid of lines within the set $EoT$; their intersection is the approximate location of the zeros as it follows from the asymptotic analysis, and they represent the vanishing of a theta function for an elliptic curve, see Section 5.2.
Consequently we are going to split the analysis in two cases;
  • the Outside of $EoT$, Sections 4.1 and 5.1;
  • the Inside of $EoT$, Sections 4.2 and 5.2.
Figure 3: From top to bottom, left to right: $(n,{{\rho}})=(16,\frac{3}{100}+\frac{13i}{100}),(16,\frac{101}{100}+\frac{13i}{100}),(17,3+1i),(17,\frac{1}{2}+\frac{i}{2}),(26,\frac{101}{100}),(40,\frac{3}{100})$. Observe that the accuracy of the localization of the zeros by the grid depends on the ratio $|{{\rho}}|/n$, the smaller the better the approximation. For example in the case in the top right $(n,{{\rho}})=(17,3+1i)$ the ratio is significant (approximately $|{{\rho}}|/n\simeq 0.19$) and an analysis where ${{\rho}}$ is considered as scaling would undoubtedly be more appropriate. In all cases we have ${K}=0$. See the explanation in Rem. 5.3.
4 Construction of the $g$–function
The method of the steepest descent analysis requires the construction of a scalar function with certain properties that we list below. This function is universally known in the literature as the $g$–function. Here we recall that the symbol of the moment functional is
$\theta(z;s)=\frac{ns}{z}+{{\rho}}\ln\left(1-\frac{1}{z}\right)+{{K}}\ln z.$ (4.1)
Since ${{K}},{{\rho}}$ will not be scaling in our setup, we introduce $\theta_{0}(z;s)=\frac{s}{z}$ (the “scaling” part) so that
@@ 4 Construction of the $g$–function
If ${{K}}=n{{K}}_{0}$ and/or ${{\rho}}=n{{\rho}}_{0}$, namely, if we were to scale ${{K}},{{\rho}}$, we would accordingly re-define $\theta_{0}$ to contain all terms proportional to $n$.
Definition 4.1 (The $g$–function and its properties)
The $g$–function is a locally bounded analytic function on ${\mathbb{C}}\setminus\Gamma$ where $\Gamma$ is a union of oriented contours (to be determined) extending to infinity satisfying the properties listed hereafter.
  1. 1.
    The contour $\Gamma$ can be written as $\Gamma=\Gamma_{m}\cup\Gamma_{c}\cup\Gamma_{\infty}$ (with $\Gamma_{m}$ denoting the “main arc(s)”, and $\Gamma_{c}$ the “complementary arc(s)”) where each of the components have pairwise disjoint relative interiors and both $\Gamma_{m},\Gamma_{c}$ consist of a finite union of compact arcs: $\Gamma_{\{m,c\}}=\bigsqcup\Gamma_{\{m,c\}}^{(j)}$. Finally $\Gamma_{\infty}$ is a simple contour extending to infinity from a finite point, traversing eventually the negative real axis and oriented from infinity.
  2. 2.
@@
  3. 3.
    For each $z\in\Gamma_{m}\cup\Gamma_{c}$ we have
    $g(z_{+})+g(z_{-})$ $=-\theta_{0}(z)-\ell+i{\varpi}_{j},\ \ \ \ \ {\varpi}_{j}\in{\mathbb{R}},\ \ \ \ z\in\Gamma_{m}^{(j)},$ (4.3)
    $g(z_{+})-g(z_{-})$ $=i{\widehat{\varpi}}_{j},\ \ \ \ \ \ {\widehat{\varpi}}_{j}\in{\mathbb{R}},\ \ \ \ z\in\Gamma_{c}^{(j)},$ (4.4)
    for some constants $\ell,\varpi_{j},\widehat{\varpi}_{j}$ (different on each of the connected components [^5] We will use different notation for these constants in the specific cases we discuss below. of $\Gamma_{m},\ \Gamma_{c}$), while
    $g(z_{+})-g(z_{-})=2i\pi,\ \ \ z\in\Gamma_{\infty}.$ (4.5)
  4. 4.
    as $z\to\infty$ in ${\mathbb{C}}\setminus\Gamma$ we have
    $g(z)=\ln(z)+\mathcal{O}(z^{-1}).$ (4.6)
  5. 5.
    The real part of the $g$–function is continuous on ${\mathbb{C}}$ (including $\Gamma$) and harmonic on ${\mathbb{C}}\setminus\Gamma_{m}$ and satisfies the following inequalities:
@@
      For all $z\in\Gamma_{c}$ we have
      $\mathrm{Re}\,\left(g(z_{+})+g(z_{-})+\theta_{0}(z)+\ell\right)=\mathrm{Re}\,\left(2g(z)+\theta_{0}(z)+\ell\right)\leq 0,$ (4.7)
@@
      For $z\in\Gamma_{m}$ we have
      $\mathrm{Re}\,\left(g(z_{+})+g(z_{-})+\theta_{0}(z)+\ell\right)=\mathrm{Re}\,\left(2g(z)+\theta_{0}(z)+\ell\right)\equiv 0.$ (4.8)
@@
      The inequality
      $\mathrm{Re}\,\left(2g(z)+\theta_{0}(z)+\ell\right)\geq 0$ (4.9)
      holds in an open neighbourhood $U$ of $\dot{\Gamma}_{m}$ (the $\dot{}$ indicating the interior set in the relative topology of the collection of arcs $\Gamma_{m}$, i.e., the arcs minus the end-points) with the equality holding only on $\Gamma_{m}$ itself.
It is convenient to reformulate Def. 4.1 in terms of the so–called effective potential
@@
  1. 1.
    $\mathrm{Re}\,(\varphi(z))$ is continuous on ${\mathbb{C}}\setminus\{0\}$, harmonic on ${\mathbb{C}}\setminus\Gamma_{m}\cup\{0\}$.
  2. 2.
    $\mathrm{Re}\,(\varphi-\theta_{0})$ is harmonic near $z=0$ and $\mathrm{Re}\,(\varphi-2\ln(z)-\theta_{0}(z))$ is harmonic at infinity.
  3. 3.
    The following equalities and inequalities hold
    $\varphi(z_{+})+\varphi(z_{-})$ $=2i{\varpi}_{j},\qquad z\in\Gamma_{m}^{(j)}$ (4.11)
    $\varphi(z_{+})-\varphi(z_{-})$ $=2i{\varsigma}_{j},\qquad z\in\Gamma_{c}^{(j)}$ (4.12)
    $\varphi(z_{+})-\varphi(z_{-})$ $=4i\pi,\qquad z\in\Gamma_{\infty}$ (4.13)
    $\mathrm{Re}\,\varphi(z)$ $\equiv 0,\ \ \ z\in\Gamma_{m}$ (4.14)
    $\mathrm{Re}\,\varphi(z)$ $\leq 0,\ \ \ z\in\Gamma_{c}$ (4.15)
    $\mathrm{Re}\,\varphi(z)$ $\geq 0,\ \ \ z\in U,$ (4.16)
    with $U$ as in Def. 4.1, 5(iii).
In the next two sections we are going to construct $\varphi$ directly and verify those properties.
4.1 Outside $EoT$: genus 00 $g$–function and effective potential
For $|s|$ sufficiently large we are going to postulate first, and then verify, the form of the effective potential. We will see that the conditions in Cor. 4.1 are fulfilled for $s$ ranging in an unbounded region and up to the boundary of $EoT$. This is the region bounded by the black arcs in Fig 3.
We start by postulating the following expression;
$\varphi(z;s):=2\int_{-\frac{is}{2}}^{z}\frac{\sqrt{w^{2}+\frac{s^{2}}{4}}}{w^{2}}{\mathrm{d}}w\ \Rightarrow\ \ \varphi^{\prime}(z;s)=2\frac{\sqrt{z^{2}+\frac{s^{2}}{4}}}{z^{2}}.$ (4.17)

### v3

@@ 3. $(3)$
$\oint_{\gamma}\frac{\widetilde{p}_{n-1}(w){\rm e}^{\theta(w)}{\mathrm{d}}w}{2{\rm i}\pi w}=-\dfrac{\begin{vmatrix}\mu_{0}&\mu_{1}&\mu_{2}&\dots&\mu_{n-1}\\ \mu_{1}&\mu_{2}&\mu_{3}&\dots&\mu_{n}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ \mu_{n-2}&\mu_{n-1}&\mu_{n}&\dots&\mu_{2n-3}\\ \mu_{-1}&\mu_{0}&\mu_{1}&\dots&\mu_{n-2}\end{vmatrix}}{\tau_{n}}$
[…]
@@
$\oint_{\gamma}\frac{p_{n}(w){\rm e}^{\theta(w)}{\mathrm{d}}w}{2{\rm i}\pi w}=\dfrac{1}{2{\rm i}\pi}\dfrac{\begin{vmatrix}\mu_{0}&\mu_{1}&\mu_{2}&\dots&\mu_{n}\\ \mu_{1}&\mu_{2}&\mu_{3}&\dots&\mu_{n+1}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ \mu_{n-1}&\mu_{n}&\mu_{n+1}&\dots&\mu_{2n-1}\\ \mu_{-1}&\mu_{0}&\mu_{1}&\dots&\mu_{n-1}\end{vmatrix}}{\tau_{n}}$
[…]
@@
Inserting (2.2), (2.2), (2.18), (2.19) into (2.17), simplifying and rearranging the rows of the resulting determinants, we obtain
@@
$\bigl(Y_{n}(0)^{-1}Y_{n}^{\prime}(0)\bigr)_{11}={}$ $-\dfrac{\begin{vmatrix}\mu_{0}&\mu_{2}&\dots&\mu_{n}\\ \mu_{1}&\mu_{3}&\dots&\mu_{n+1}\\ \vdots&\vdots&\ddots&\vdots\\ \mu_{n-1}&\mu_{n+1}&\dots&\mu_{2n-1}\end{vmatrix}}{\tau_{n}}\dfrac{\begin{vmatrix}\mu_{-1}&\mu_{0}&\mu_{1}&\dots&\mu_{n-2}\\ \mu_{0}&\mu_{1}&\mu_{2}&\dots&\mu_{n-1}\\ \mu_{1}&\mu_{2}&\mu_{3}&\dots&\mu_{n}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ \mu_{n-2}&\mu_{n-1}&\mu_{n}&\dots&\mu_{2n-3}\end{vmatrix}}{\tau_{n}}$
$+\dfrac{\begin{vmatrix}\mu_{0}&\mu_{2}&\dots&\mu_{n-1}\\ \mu_{1}&\mu_{3}&\dots&\mu_{n}\\ \vdots&\vdots&\ddots&\vdots\\ \mu_{n-2}&\mu_{n}&\dots&\mu_{2n-3}\end{vmatrix}}{\tau_{n}}\dfrac{\begin{vmatrix}\mu_{-1}&\mu_{0}&\mu_{1}&\dots&\mu_{n-1}\\ \mu_{0}&\mu_{1}&\mu_{2}&\dots&\mu_{n}\\ \mu_{1}&\mu_{2}&\mu_{3}&\dots&\mu_{n+1}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ \mu_{n-1}&\mu_{n}&\mu_{n+1}&\dots&\mu_{2n-1}\end{vmatrix}}{\tau_{n}}.$
@@
We now use the Desnanot–Jacobi identity, which can be stated as follows. Given a square matrix $M$ denote by $M^{[a_{1},\dots,a_{r}][b_{1},\dots,b_{r}]}$ the matrix obtained by deleting the rows $a_{1},\dots,a_{r}$ and columns $b_{1},\dots,b_{r}$}. Then
@@
$M:=\begin{vmatrix}\mu_{-1}&\mu_{0}&\mu_{1}&\dots&\mu_{n-1}\\ \mu_{0}&\mu_{1}&\mu_{2}&\dots&\mu_{n}\\ \mu_{1}&\mu_{2}&\mu_{3}&\dots&\mu_{n+1}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ \mu_{n-1}&\mu_{n}&\mu_{n+1}&\dots&\mu_{2n-1}\end{vmatrix}\in\mathfrak{gl}(n+1).$
@@
$\partial_{t}\mu_{j}(t)=\oint_{\gamma}z^{j-1}\left(1-\frac{1}{z}\right)^{{\rho}}z^{{K}}{\rm e}^{\frac{t}{z}}{\mathrm{d}}z=\mu_{j-1}(t),\qquad j\in{\mathbb{Z}}.$
This allows us to show that $\partial_{t}\det\tau_{n}=\det M^{[n+1][2]}$ and hence $\bigl(Y_{n}(0)^{-1}Y_{n}^{\prime}(0)\bigr)_{11}=-\partial_{t}\ln\tau_{n}(t)$. This concludes the proof. ∎
@@ 3 Asymptotic analysis
The goal of this second part of the paper is to study the behaviour of the poles of the rational solution or, which is the same, the zeros of the Hankel determinant $\tau_{n}(t)$ as $n\to\infty$. More precisely, and in line with similar investigations done for rational solution of PII [6, 17, 18], we are going to re-scale the zeros concurrently with $n$ and study the behaviour of the zeros in the $s=\frac{t}{n}$-plane. Namely, we study the sequence of functions (polynomials)
$\mathcal{T}_{n}(s):=\tau_{n}(ns;{K},{{\rho}}),$ (3.1)
@@
$\tau_{n}(t;{K},{{\rho}}):=\det\left[\oint_{\gamma}z^{a+b+{K}}\left(1-\frac{1}{z}\right)^{{\rho}}{\rm e}^{\frac{t}{z}}\frac{{\mathrm{d}}z}{2{\rm i}\pi}\right]_{a,b=0}^{n-1}.$
We emphasize that we are not scaling the parameters $\rho$$\in{\mathbb{C}}\setminus{\mathbb{Z}}$ nor $K\in{\mathbb{Z}}$ as $n\to\infty$. A different asymptotic analysis would be needed if we were to set ${{\rho}}\propto n$ and/or $K\propto n$. We will briefly point out the main differences when appropriate.
Theorem 2.5 can be re-formulated for convenience in the following form.
Theorem 3.1.
The function $\mathcal{T}_{n}(s):=\tau_{n}(ns;K,{{\rho}})$ (3.1) is zero if and only if the Riemann–Hilbert Problem 2.4 has no solution ((with $t=ns)$.
The asymptotic analysis of Riemann–Hilbert Problem 2.4 with $t=ns$ falls within the purview of the Deift–Zhou asymptotic method that was brought to fruition in [26] for orthogonal polynomials, see also [11, 25].
@@
  1. (1)
    Construct a “$g$-function”. This is a function $g(z;s)$ whose domain of analyticity $\mathscr{D}_{s}$ depends (continuously) on the parameter $s$. It can be expressed as an Abelian integral on a hyperelliptic Riemann surface $\mathcal{R}$ whose genus depends on $s$. The $g$-function is uniquely characterized by a free-boundary value problem and certain inequalities which will be described in Section 4.
  2. (2)
    Use the constructed $g$-function to normalize Riemann–Hilbert Problem 2.4 and express a new matrix-valued function $W(z)$ which satisfies a new RHP.
  3. (3)
@@
The key concept is that the solvability of the initial Riemann–Hilbert Problem 2.4 is equivalent to the solvability of its final reformulation hinted at in the list above. Thus, if we can guarantee that the final reformulation is solvable for $s$ in a suitable domain, we can conclude that there are no zeros of $\mathcal{T}_{n}(s)$ in that domain (at least for $n$ sufficiently large).
Let us give a visual overview of the result. In Figure 1, we can see several instances of plots of the zeros of $\mathcal{T}_{n}(s)$ for various values of $\rho$}}. The common feature, which will be proved, is that all the zeros lie, asymptotically for large $n$, within the “Eye of the Tiger” region ($\rm EoT$}), marked by the black arcs. The equation of these black arcs is given implicitly in (4.7).
Also indicated is a grid of lines within the set $\rm EoT$}. Their intersection is the approximate location of the zeros as it follows from the asymptotic analysis, and they represent the vanishing of a theta function for an elliptic curve, see Section 5.2.
Consequently we are going to split the analysis in two cases:
  • the outside of $\rm EoT$}, Sections 4.1 and 5.1,
  • the inside of $\rm EoT$}, Sections 4.2 and 5.2.
4 Construction of the $\boldsymbol{g}$-function
The method of the steepest descent analysis requires the construction of a scalar function with certain properties that we list below. This function is universally known in the literature as the $g$-function. Here we recall that the symbol of the moment functional is
$\theta(z;s)=\frac{ns}{z}+{{\rho}}\ln\left(1-\frac{1}{z}\right)+{{K}}\ln z.$
Since K{{K}}, $\rho$}} will not be scaling in our setup, we introduce $\theta_{0}(z;s)=\frac{s}{z}$ (the “scaling” part) so that
@@ 4 Construction of the $\boldsymbol{g}$-function
If ${{K}}=n{{K}}_{0}$ and/or ${{\rho}}=n{{\rho}}_{0}$, namely, if we were to scale K{{K}}, $\rho$}}, we would accordingly re-define $\theta_{0}$ to contain all terms proportional to $n$.
Definition 4.1 (the $g$-function and its properties).
The $g$-function is a locally bounded analytic function on ${\mathbb{C}}\setminus\Gamma$ where $\Gamma$ is a union of oriented contours (to be determined) extending to infinity satisfying the properties listed hereafter.
  1. (1)
    The contour $\Gamma$ can be written as $\Gamma=\Gamma_{m}\cup\Gamma_{c}\cup\Gamma_{\infty}$ (with $\Gamma_{m}$ denoting the “main arc(s)”, and $\Gamma_{c}$ the “complementary arc(s)”) where each of the components have pairwise disjoint relative interiors and both $\Gamma_{m}$, $\Gamma_{c}$ consist of a finite union of compact arcs: $\Gamma_{\{m,c\}}=\bigsqcup\Gamma_{\{m,c\}}^{(j)}$. Finally, $\Gamma_{\infty}$ is a simple contour extending to infinity from a finite point, traversing eventually the negative real axis and oriented from infinity.
  2. (2)
@@
  3. (3)
    For each $z\in\Gamma_{m}\cup\Gamma_{c}$, we have
    $g(z_{+})+g(z_{-})=-\theta_{0}(z)-\ell+{\rm i}{\varpi}_{j},\qquad{\varpi}_{j}\in{\mathbb{R}},\qquad z\in\Gamma_{m}^{(j)},$
    $g(z_{+})-g(z_{-})={\rm i}{\widehat{\varpi}}_{j},\qquad{\widehat{\varpi}}_{j}\in{\mathbb{R}},\qquad z\in\Gamma_{c}^{(j)},$
    for some constants $\ell$, $\varpi_{j}$, $\widehat{\varpi}_{j}$ (different on each of the connected components [^3] We will use different notation for these constants in the specific cases we discuss below. of $\Gamma_{m}$, $\Gamma_{c}$), while $g(z_{+})-g(z_{-})=2{\rm i}\pi$, $z\in\Gamma_{\infty}$.
  4. (4)
    As $z\to\infty$ in ${\mathbb{C}}\setminus\Gamma$, we have $g(z)=\ln(z)+\mathcal{O}\bigl(z^{-1}\bigr)$.
  5. (5)
    The real part of the $g$-function is continuous on $\mathbb{C}$} (including $\Gamma$) and harmonic on ${\mathbb{C}}\setminus\Gamma_{m}$ and satisfies the following inequalities:
@@
      For all $z\in\Gamma_{c}$, we have
      $\operatorname{Re}\left(g(z_{+})+g(z_{-})+\theta_{0}(z)+\ell\right)=\operatorname{Re}\left(2g(z)+\theta_{0}(z)+\ell\right)\leq 0,$
@@
      For $z\in\Gamma_{m}$, we have
      $\operatorname{Re}\left(g(z_{+})+g(z_{-})+\theta_{0}(z)+\ell\right)=\operatorname{Re}\left(2g(z)+\theta_{0}(z)+\ell\right)\equiv 0.$
@@
[…]
It is convenient to reformulate Definition 4.1 in terms of the so-called effective potential
@@
  1. $(1)$
    $\operatorname{Re}(\varphi(z))$ is continuous on ${\mathbb{C}}\setminus\{0\}$, harmonic on ${\mathbb{C}}\setminus\Gamma_{m}\cup\{0\}$.
  2. $(2)$
    $\operatorname{Re}(\varphi-\theta_{0})$ is harmonic near $z=0$ and $\operatorname{Re}(\varphi-2\ln(z)-\theta_{0}(z))$ is harmonic at infinity.
  3. $(3)$
    The following equalities and inequalities hold:
    $\varphi(z_{+})+\varphi(z_{-})=2{\rm i}{\varpi}_{j},\qquad z\in\Gamma_{m}^{(j)},\qquad\varphi(z_{+})-\varphi(z_{-})=2{\rm i}{\varsigma}_{j},\qquad z\in\Gamma_{c}^{(j)},$
    $\varphi(z_{+})-\varphi(z_{-})=4{\rm i}\pi,\qquad z\in\Gamma_{\infty},\qquad\operatorname{Re}\varphi(z)\equiv 0,\qquad z\in\Gamma_{m},$
    $\operatorname{Re}\varphi(z)\leq 0,\qquad z\in\Gamma_{c},\qquad\operatorname{Re}\varphi(z)\geq 0,\qquad z\in U,$
    with $U$ as in Definition 4.1(5)(iii)(5)\,(iii).
In the next two sections, we are going to construct $\varphi$ directly and verify those properties.
4.1 Outside $\boldsymbol{{\rm EoT}}$: Genus $\boldsymbol{0}$ $\boldsymbol{g}$-function and effective potential
For $|s|$ sufficiently large, we are going to postulate first, and then verify, the form of the effective potential. We will see that the conditions in Corollary 4.2 are fulfilled for $s$ ranging in an unbounded region and up to the boundary of $\rm EoT$}. This is the region bounded by the black arcs in Figure 3.
Figure 3: From top to bottom, left to right: $(n,{{\rho}})=\big(16,\frac{3}{100}+\frac{13{\rm i}}{100}\big),\big(16,\frac{101}{100}+\frac{13{\rm i}}{100}\big),(17,3+1{\rm i}\big),\allowbreak\big(17,{\frac{1}{2}+\frac{\rm i}{2}}\big),\big(26,\frac{101}{100}\big),\big(40,\frac{3}{100}\big)$. Observe that the accuracy of the localization of the zeros by the grid depends on the ratio $|{{\rho}}|/n$, the smaller the better the approximation. For example, in the case in the top right $(n,{{\rho}})=(17,3+1{\rm i})$ the ratio is significant (approximately $|{{\rho}}|/n\simeq 0.19$) and an analysis where $\rho$}} is considered as scaling would undoubtedly be more appropriate. In all cases, we have K=0{K}=0. See the explanation in Remark 5.19.
We start by postulating the following expression:
$\varphi(z;s):=2\int_{-\frac{{\rm i}s}{2}}^{z}\frac{\sqrt{w^{2}+\frac{s^{2}}{4}}}{w^{2}}{\mathrm{d}}w\ \Rightarrow\ \ \varphi^{\prime}(z;s)=2\frac{\sqrt{z^{2}+\frac{s^{2}}{4}}}{z^{2}}.$ (4.3)